OpenStudy (anonymous):

A large container has the shape of a frustum of a cone with top radius 8m, bottom radius 2m, and height 6m. The container is being filled with water at the constant rate of 2.9m^3/min. At what rate is the level of water rising at the instant the water is 1m deep? Answer: 10.3cm/min

OpenStudy (anonymous):

\[\frac{ dV }{ dt } = +2.9, \frac{ dH }{ dt }=?\]

OpenStudy (anonymous):

@iop360 Do you have the Volume formula?

OpenStudy (anonymous):

of a regular cone, yes

OpenStudy (anonymous):

\[V = \frac{ 1 }{ 3 }(\pi)r^2h\]

OpenStudy (anonymous):

No this one, have 2 bases!

OpenStudy (anonymous):

i think what we are supposed to do though is make it into two cones though

OpenStudy (anonymous):



OpenStudy (anonymous):

R is top radius?

OpenStudy (anonymous):

b and B represent the different volumes?

OpenStudy (anonymous):

how did you derive this formula

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ill try it. thanks!

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

have you tried it?

OpenStudy (anonymous):

im getting a wrong answer

OpenStudy (anonymous):

i got .383...m/min

OpenStudy (anonymous):

answer is supposed to be 10.3 cm/min

OpenStudy (anonymous):

oh wait i think i made an error.. let me recalculate

OpenStudy (anonymous):

my expression for V is: \[V = [\frac{ 16 }{ 27 }(\pi)h^3 + \frac{ 4}{ 27 }(\pi)h^3 + \frac{ 1 }{ 9 }(\pi)h^3]\]

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

oh wait.. 1/27, not 1/9

OpenStudy (anonymous):

ok yeah thats what i get now

OpenStudy (anonymous):

taking its derivative, you get \[\frac{ 7 }{ 3 }(\pi)h^2\]

OpenStudy (anonymous):

dh/dt beside it of course

OpenStudy (anonymous):

(2.9)/(7pi/3) = dh/dt after you plug h =1, which does nothing

OpenStudy (anonymous):

0.395...m^3/min

OpenStudy (anonymous):

i think we did it right, not sure why its resulting in the wrong answer

OpenStudy (anonymous):

did you get 0.395...m/min

OpenStudy (anonymous):

doesnt that still result in 39.5cm/min then?

OpenStudy (anonymous):

it might be a solution, but i dont get it

OpenStudy (anonymous):

hmm is there anything extra in the formula you may have forgotten

OpenStudy (anonymous):

http://www.youtube.com/watch?v=1v1Pp-lJSKY this video ?

OpenStudy (anonymous):

http://jwilson.coe.uga.edu/emt725/Frustum/Frustum.cone.html this link doesnt have a square root in the formula

OpenStudy (anonymous):

bleh, the formula without the sq.root doesnt work either...

OpenStudy (anonymous):

!!! i think i got it

OpenStudy (anonymous):

forgot the pi from the above forumla

OpenStudy (anonymous):

i got 10.1 though...calculator didnt take exact values

OpenStudy (anonymous):

\[V = \frac{ (\pi)h }{ 3 }[R^2 + Rr + r^2]\]

OpenStudy (anonymous):

im going to calculate it again..

OpenStudy (anonymous):

im getting 10.05... cm/min that is slightly off

OpenStudy (anonymous):

hm ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im going to try it with different values

OpenStudy (anonymous):

oh wait...i used base for r in that last formula

OpenStudy (anonymous):

uh oh

OpenStudy (anonymous):

ill try it with r/R

OpenStudy (anonymous):

annnnnnd were back to 21pi/27 h^3

OpenStudy (anonymous):

yeah, i accidently used the B/b for that forumla instead of R/r

OpenStudy (anonymous):

and yet i got close to a right answer

OpenStudy (anonymous):

do we sub in h = 1 or h =6? h = 1 is right isnt it?

OpenStudy (anonymous):

Yes, we calculate the instant rate when h = 1

OpenStudy (anonymous):

the other guy reading this question: are you doing this question?

OpenStudy (callisto):

*learning* ._.!

OpenStudy (anonymous):

i could switch the values for the question and try it again if you want, haha

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

a large container has the shape of a frustum of a cone with top radius 10m, bottom radius 6m, and height 4m. The container is being filled with water at the constant rate of 2.9m^3/min. At what rate is the level of water rising at the instant rate the water is 2m deep?

OpenStudy (anonymous):

answer: 1.4cm/min

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i get 1.9 ..hmmm

OpenStudy (anonymous):

it could be a glitch with the thing, maybe.

OpenStudy (anonymous):

ok, im pretty sure were doing it right. imma close this down

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

Check with your teacher about the formula! I'll sure will message you if I find out something interesting :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I'm so sorry that I'm underestimate this Related Rate frustum of Cone! Here's the correct logic: V = π/ 3 ( R² + Rr + r² ) H With r = 2 and R is Radius variable and H is the Height variable -> V = π/ 3 ( R² + 2R+ 4 ) H ( 1 ) From the ratio of right triangle: ( R - 2 ) / ( 8 - 2 ) = H/ 6 --> R = H + 2 (2) Plug (2) into (1): => V = π/ 3 [ ( H +2) + 2( H +2) + 4 ] H = π/ 3 [ H³ + 6H² + 12H ] so V' = π/ 3 [ 3H² + 12H + 12 ] H' At H = 1: V' = 27 π/ 3 * H' 2.9 = 9 π * H' Thus H' = 2.9 / 9π = .1025 m/ min = 10.3 cm/ min

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@Chlorophyll one question though.. what do you do exactly during the ratios of the triangle part? eg. say my top radius was 9, bottom was 8, and height was 5 how do you set it up exactly?

OpenStudy (anonymous):

Ratio R and H: ( R - 8 )/ ( 9-8) = H/5 -> R = ( H/5 ) + 8